Sketch the region enclosed by the given curves. Then, find its area by integrating with respect to y. x=4-y’, x= y² - 4
Answer #1

fagh x=A-ya x= y²4 y=2 x=442 2 4-12 x=yzy area = SS dxdy yer x=yly y=-2 2 area ļ [ (4-y?)-(y²4) Jay - 2ye] dy 2 fuention 2 (8